OpenStudy (anonymous):

A dance CD has 13 songs on it - 7 are slow, and 6 are fast. When the DJ at a dance plays a song it is not played again. All songs from the CD are played at random. What is the probability that the first two songs played are slow songs? a. 49/156 b. 42/156 c. 42/169 d. 49/169

OpenStudy (anonymous):

\[\frac{7}{13}\times \frac{6}{12}\] assuming he the dj does not repeat the same song twice

OpenStudy (anonymous):

Okay, thank you.